# AstroLoc: Robust Space to Ground Image Localizer

### v1

@@ Abstract
Astronauts take thousands of photos of Earth per day from the International Space Station, which, once localized on Earth’s surface, are used for a multitude of tasks, ranging from climate change research to disaster management. The localization process, which has been performed manually for decades, has recently been approached through image retrieval solutions: given an astronaut photo, find its most similar match among a large database of geo-tagged satellite images, in a task called Astronaut Photography Localization (APL). Yet, existing APL approaches are trained only using satellite images, without taking advantage of the millions open-source astronaut photos. In this work we present the first APL pipeline capable of leveraging astronaut photos for training. We first produce full localization information for 300,000 manually weakly labeled astronaut photos through an automated pipeline, and then use these images to train a model, called AstroLoc. AstroLoc learns a robust representation of Earth’s surface features through two losses: astronaut photos paired with their matching satellite counterparts in a pairwise loss, and a second loss on clusters of satellite imagery weighted by their relevance to astronaut photography via unsupervised mining. We find that AstroLoc achieves a staggering 35% average improvement in recall@1 over previous SOTA, pushing the limits of existing datasets with a recall@100 consistently over 99%. Finally, we note that AstroLoc, without any fine-tuning, provides excellent results for related tasks like the lost-in-space satellite problem and historical space imagery localization.
@@ 1 Introduction
Figure 1: One model, many space to ground applications. We train a single model, AstroLoc, that succeeds in multiple space-based image retrieval settings with significant domain gaps. We show strong performance on astronaut photography localization, “lost in space” orbit determination, and historical (Space Shuttle) photography localization. Each group of 3 images represent a query and its top-2 predictions from searching over a worldwide database of thousands of satellite images. Correct predictions in green, wrong predictions in red.
@@
Earth observation is the process of collecting information about the Earth’s surface, which is vital for monitoring the state of our planet. Among the multitude of systems used to acquire Earth observations data, most of which rely on automatic collection by nadir-facing satellites, manual acquisition of photos by astronauts aboard the International Space Station (ISS) is a clear outlier. This imagery has distinctive properties including (1) a wide range of spatial resolutions (up to 2 meters per pixel), (2) the potential for oblique perspectives to observe topography and height (e.g., measure the height of a volcanic eruption to divert flights), (3) various illumination conditions (whereas satellite imagery is typically sun-synchronous, always capturing the same area at the same time of day), and (4) offering the highest resolution open source Earth observations data. Furthermore, due to the ISS’s 90 minute repeating orbit, astronauts can be tasked with imaging for near real time disaster response, and they can use human intuition to focus on relevant areas and events, making the photographs very information-dense, whereas satellites commonly collect images of areas regardless of semantic value. All these characteristics make this data source a perfect complement to satellite data and a crucial component for climate science [23], atmospheric phenomena [15, 36], urban planning [8, 13, 29, 26], and, most importantly, disaster management and response [30] [^1] https://storymaps.arcgis.com/stories/947eb734e811465cb0425947b16b62b3.
@@
Although over 5 million photos have been collected from the ISS since the beginning of its operations, unfortunately their full potential has yet to be unlocked. The complication is that astronaut photos, unlike satellite imagery, are not automatically localized w.r.t. the Earth’s surface. Even though the position of the ISS can be determined from the timestamp of the photo and the ISS orbit path, an astronaut can point their camera (a standard hand-held DSLR) toward any location on Earth within their vast visibility range, which spans roughly 20 million square kilometers (sqkm) [^2] Considering that the ISS orbits at an altitude of 415 kilometers.. Therefore, localizing a single photo that can cover an area as small as 100 sqkm (i.e. 0.0005% of the visible/search area) is akin to finding a needle in a haystack. For example, when an astronaut is above Texas and photographs a wildfire they see, the burning area could be anywhere from Canada to Mexico depending on the camera orientation, and the only way to know where is to geolocate the image.
@@
The potential value these photos hold when geolocated has led to a concerted manual localization effort, with experts and citizen scientists localizing over 300,000 of these images – a process defined by NASA as a “monumentally important, but monumentally time-consuming job” [^3] We highly encourage the reader to view this clip https://www.youtube.com/watch?v=drrP_Iss0gA&t=295s. Nevertheless, manually geolocating a single astronaut photo can take hours, effectively making this procedure inadequate given the ever-growing number of images that are being collected (up to 10k per day). This problem has prompted recent investigation into automating the task of Astronaut Photography Localization (APL) with deep learning models, combining both image retrieval [5] and image matching techniques [32, 4, 33]. Image retrieval is used as a first step, to search within a database of precisely geo-referenced satellite images for those that are most similar to the query astronaut photo. With such a system, a query can be coarsely localized in less than a second, making it a viable solution for such a large scale problem. Afterwards, the retrieval candidates may be refined using more computationally demanding matching techniques. Although these pipelines have shown satisfactory speeds for APL, their performance show a wide margin for improvement, and we hypothesize that this is due to existing APL models being trained with only satellite images [5]. This is a critical limitation, considering that the end goal is to geo-reference astronaut photos, not satellite images.
@@
We argue that it would be beneficial to exploit the available 300,000 hand-labeled open-source astronaut photos to train the image retrieval model. The difficulty with doing so is that these photos are only weakly annotated, i.e., the manual label only provides the GPS location of a random point close to the center of the image. Therefore, these weakly labeled photos cannot be directly used in a metric learning framework, where the model must be provided with pairs of queries (astronaut photos) and database (satellite) images from the same location. In order to meet this criteria, we need these photos to be labeled with their full footprint (i.e. precise GPS location of its four corners), so that we can exactly determine images from the database to be used as positive samples (i.e., with a good overlap w.r.t. the training queries). The concept of weak and full labels is visualized in Fig. 2.
@@
Figure 2: Visual example of weak (manual) annotation and full annotation of an astronaut photo. Weak annotation is the GPS coordinates of a single point, which does not provide information regarding the image’s size (i.e. it could cover a city or an entire continent), whereas full annotation provides GPS for its 4 corners (called the footprint), from which the coordinate of any pixel within the image can be easily calculated (pixel-wise label).
@@
In light of these considerations, we first produce a precise annotation of these 300,000 photos, leveraging their weak annotation and a state-of-the-art matching pipeline to label their full footprint. Then, we demonstrate the importance of using this newly annotated data for training, by extending the state-of-the-art method EarthLoc [5] in two ways. First, we create pairs of matching astronaut-satellite images, and use them in a pairwise contrastive loss that takes advantage of the fact that the two images within each pair come from different domains. The creation of these pairs relies on the new annotation, i.e., we select a pair if the two images have enough intersection over union on Earth’s surface. Even with this loss that takes advantage of the newly labeled 300,000 astronaut photos, it is still important to fully leverage the much larger set of 5.3M geo-referenced satellite images, as is done in previous methods. However, we observe that the satellite images are uniformly distributed on the surface of the Earth, whereas the photos from astronauts tend to be unevenly distributed, with higher concentrations in salient areas like volcanoes, glaciers, lakes and coastlines. Therefore, we use a second contrastive loss that leverages tuples of satellite images only, but sampled according to the distribution of astronaut photos.
We find that our training pipeline, yielding a model we call AstroLoc, leads to a model so powerful that it saturates existing test sets (recall@100 above 99%), and we therefore propose new, more challenging test sets that better reflect the task of APL in the real world.
To summarize, our contributions are as follows:
  • we provide precise localization labels to 300,000 weakly labeled astronaut photographs, which will be made available to support further research in this field;
  • we propose two techniques to effectively adopt this data source into an Astronaut Photography Localization training pipeline;
  • AstroLoc leads to a staggering 35% improvement on average of recall@1 over previous SOTA, and practically solves existing test sets (recall@100 around 99%);
@@
Furthermore, AstroLoc is practically useful and has already been used to localize hundreds of thousand of images available here [^4] Photos available at: https://eol.jsc.nasa.gov/ExplorePhotos/ , and shows great potential for tasks like the lost-in-space satellite and historical space image localization. We finally note that, thanks to AstroLoc, in a few months the backlog of non-localized astronaut photos will be nearly empty for the first time since the launch of the ISS, with the exception of a small fraction that still produce failure cases.
@@ 2 Related Work
While these solutions have advanced the field of Earth from space localization, they are limited, as Berton et al. [5] notes. The matching-based methods can not achieve the scalability and speed required to localize astronaut photographs in real time. Berton et al. proposes approaching the task as an image retrieval problem, where the database is composed of satellite images of known location. In this setting, AnyLoc [17], a universal place recognition model, has been shown to produce strong performance and robustness to the query-database domain gap.
Unmanned aerial vehicle localization A closely related task is that of unmanned aerial vehicle (UAV) localization. Similar to Astronaut Photography Localization (APL), it is approached through cross-domain image retrieval, where the database is often made of satellite imagery and the queries are photos collected by a drone. The goal is to efficiently find the best prediction for a given query in order to localize the drone in real time [38]. Among notable examples from the literature, Bianchi and Barfoot [6] introduced a method utilizing autoencoded satellite images, significantly reducing storage and computation costs while maintaining robust localization performance. Dai et al. [7] proposed an end-to-end framework, FPI, that directly identifies UAV locations by matching UAV-view images with satellite-view counterparts, streamlining the localization process. Li et al. [18] developed a transformer-based adaptive semantic aggregation method, focusing on part-level representations to improve UAV-satellite image matching by capturing detailed semantic information.
@@ 3 Data
Our goal is to localize the extent of astronaut photos (queries) given a worldwide map of geo-referenced satellite images (database). Astronaut photographs represent one of the most diverse and long standing Earth observations data sources. The dataset contains a wide range in spatial resolution, field of view, illumination (including night imagery), and obliquity (astronauts can tilt the camera and take oblique photos). These images have many applications in numerous fields [15, 36, 23, 8, 13, 29, 26], and most importantly in disaster management and response [30] [^5] https://eol.jsc.nasa.gov/Collections/Disasters/ShowDisastersCollection.pl.
For database images we follow [5] and use a composited set of cloudless open-source satellite yearly imagery from Sentinel-2, named S2. We use tiles from zoom levels 8-12 [24] (expanding in both directions on the set used in [5]), to provide a thorough train and test set that reflects the extents of the query images. Characteristics of astronaut and satellite imagery are compared in Tab. 1.
@@ 3.1 Training Set
In contrast to previous work, we propose taking advantage of the 300k manually localized astronaut photos for training. The idea is to pair these photos with matching satellite imagery from the database (i.e. with enough Intersection over Union, or IoU) and apply contrastive training to train models that are robust to the domain differences between astronaut and satellite imagery. Unfortunately, the available manual localization is only a weak annotation (see Fig. 2), which does not provide enough information to compute IoU. To overcome this limitation, we estimate the precise footprint for each query image. For each manually localized query, we select as potential positives all the database images which contain the weak label point at each zoom level, and we rotate these potential positives by 90°, 180° and 270° to maximize the chance of finding a similar image among the potential positives. In practice, given five zoom levels, four rotations, and the fact that each point is covered by four images (there is overlap among the tiles), there are $5\times 4\times 4=80$ potential positives per query. Note that different zoom levels are required because it is impossible to estimate an image extent with only a weak label, and a photo could cover an area as small as a city or as big as a continent. We then perform image matching with SuperPoint [9] + LightGlue [19] and the EarthMatch pipeline [4] to get the footprint coordinates of each query.
This localization was successful for 221k queries, with the remaining ones not localizable due to either errors in manual labeling, too much obstruction from clouds, or the image being of the horizon (i.e. Earth limb photograph, with two corners representing outer space, hence an invalid footprint). With this now precisely localized collection of astronaut photography, we pair each query with the all database images with an IoU over $t_{iou}=0.2$ producing 865k query-database training pairs. We then discard any pair containing a query that is included within the test sets to avoid any data leakage.
@@ 3.2 Evaluation Sets
We therefore propose new evaluation sets, to include all available geolocated queries within an evaluation area, as well as add zoom levels 8 and 12 to the database to match the wider range of query areas. To avoid introducing complexity, we use the same regions as the test sets from [5]. These datasets were named after the geographic location of their center, like Texas, Gobi, and Amazon, so we call these new extended versions Texas-L (L for Large), Gobi-L, etc.
@@ 4.1 Setting
Figure 4: Overview of AstroLoc training pipeline. The upper branch feeds pairs of matching query-database images to the pairwise loss (Sec. 4.2). The lower branch (Sec. 4.3) first creates clusters of satellite images, then queries are assigned to these clusters, and then the clusters are sampled according to how many queries are assigned to each cluster (unsupervised mining). A training batch (mined quadruplets, i.e. tuples of 4 images) is sampled from a single cluster, and fed to the Multi-similarity Unsupervised Mining (MUM) loss. Queries are not fed to the MUM loss, and are not used to compute the clusters — they are only used to sample training data from a closer distribution to the queries’.
Task Astronaut Photography Localization (APL) is the task of estimating the geographic location covered by a photo (query) taken by astronauts. APL can be approached through image retrieval: for each query, estimate its location by finding the most similar match in a large database of geolocated images of Earth.
Formalization We pose the task as a retrieval problem, and rely on three sets of images:
@@
  • The test queries $\mathcal{T}=\{t_{1},t_{2},\ldots,t_{N_{\mathcal{T}}}\}$ is another set of $N_{\mathcal{T}}$ astronaut photos, disjoint from $\mathcal{Q}$, each characterized by its RGB content and timestamp. The timestamp is crucial in determining the location of the ISS when the photo was taken, which enables narrowing down the search space from a world-wide area (510M sqkm) to the local area visible from the ISS at that time (20M sqkm). Note that due to cosmic radiation bit flips, the timestamp can sometimes be incorrect: in such cases, it is necessary to conduct a “world-wide” search (see Sec. 5.5).
Training with Astronaut Photos (Queries) While various approaches have been used for APL, no previous work has taken advantage of the huge dataset of 300k astronaut photos labeled by human experts. We argue that their use at training time can lead to huge improvements in performance, and we present two novel techniques to use this data. In Sec. 4.2 we show how we use pairs of matching query-reference images (i.e., depicting the same location) described in Sec. 3.1 to implement a pairwise loss for this setup (see Fig. 4 top). Additionally, we also use a second loss computed on tuples of satellite images, by sampling them using the distribution of the queries (see an overview in Fig. 4 bottom), allowing us to use the entire set of database images for training (whereas the pairwise loss can be applied only on areas covered by astronaut photos) while still leveraging the queries, as we explain in Sec. 4.3.

### v2

@@ Abstract
Thousands of photos of Earth are taken every day by astronauts from the International Space Station. Localizing these photos, which has been performed manually for decades, has recently been approached through image retrieval solutions: given an astronaut photo, the goal is to find its most similar match among a large database of geo-tagged satellite images, in a task called Astronaut Photography Localization (APL). Yet, existing APL approaches are trained only using satellite images, without taking advantage of the millions of open-source astronaut photos. In this work we present the first APL pipeline capable of leveraging astronaut photos for training. We first produce full localization information for 300,000 manually weakly-labeled astronaut photos through an automated pipeline, and then use these images to train a model, called AstroLoc. AstroLoc learns a robust representation of Earth’s surface features through two objective functions: pairing astronaut photos with their matching satellite counterparts in a pairwise loss, and a second loss on clusters of satellite imagery weighted by their relevance to astronaut photography through unsupervised mining. AstroLoc achieves a staggering 35% average improvement in recall@1 over previous SOTA, reaching a recall@100 consistently over 99% for existing datasets. Moreover, without fine-tuning, AstroLoc provides excellent results for related tasks like the lost-in-space satellite problem and historical space imagery localization.
@@ 1 Introduction
Figure 1: One model, many space to ground applications. We train a single model, AstroLoc, that succeeds in multiple space-based image retrieval settings: astronaut photography localization, “lost in space” orbit determination, and historical (Space Shuttle) photography localization. In this figure, each group of 3 images represent a query and its top-2 predictions from searching over a worldwide database of millions of satellite images. Correct predictions in green, wrong predictions in red.
footnotetext: *Work was done outside of Amazon
@@
Earth observation is the process of collecting information about the Earth’s surface, which is vital for monitoring the state of our planet. Among the multitude of systems used to acquire Earth observations data, most of which rely on automatic collection by nadir-facing (straight-down) satellites, manual acquisition of photos by astronauts aboard the International Space Station (ISS) is a clear outlier. This imagery has distinctive properties including (1) a wide range of spatial resolutions (up to 2 meters per pixel), (2) the potential for oblique perspectives to observe topography and height (e.g., measure the height of a volcanic eruption to divert flights), (3) various illumination conditions (whereas satellite imagery is typically sun-synchronous, always capturing the same area at the same time of day), and (4) offers the highest resolution open source Earth observations data. Furthermore, due to the ISS’s 90 minute repeating orbit, astronauts can be tasked with imaging for near real time disaster response, and they can use human intuition to focus on relevant areas and events, making the photographs very information-dense, whereas satellites commonly collect images systematically, regardless of semantic value. All these characteristics make this data source a perfect complement to satellite data and a crucial component for climate science [26], atmospheric phenomena [18, 40], urban planning [9, 14, 32, 29], and, most importantly, disaster management and response [33] [^1] https://storymaps.arcgis.com/stories/947eb734e811465cb0425947b16b62b3.
@@
Although over 5 million photos have been collected from the ISS since the beginning of its operations, unfortunately their full potential has yet to be unlocked. The complication is that astronaut photos, unlike satellite imagery, are not automatically geolocated w.r.t. the Earth’s surface. Even though the position of the ISS can be determined from the timestamp of the photo and the ISS orbit path, an astronaut can point their camera (a standard hand-held DSLR) toward any location on Earth within their vast visibility range, which spans roughly 20 million square kilometers (sqkm) [^2] Considering that the ISS orbits at an altitude of 415 kilometers.. Therefore, localizing a photo that covers an area of 100 sqkm (i.e. 0.0005% of the visible/search area) is like finding a needle in a haystack. For example, when an astronaut is orbiting above Texas and photographs a wildfire, the burning area could be anywhere from Canada to Mexico depending on camera orientation, and the only way to know where is by geolocating the image.
@@
The potential value these photos hold when geolocated has led to a concerted manual localization effort, with experts and citizen scientists localizing over 300,000 of these images – a process defined by NASA as a “monumentally important, but monumentally time-consuming job” [^3] We highly encourage the reader to view this clip https://www.youtube.com/watch?v=drrP_Iss0gA&t=295s. Nevertheless, manually geolocating a single astronaut photo can take hours (hundreds of thousands of human-hours have been spent on hand-labeling 300k of them), effectively making this procedure inadequate given the ever-growing number of images that are being collected (up to 10k per day). This problem has prompted recent investigation into automating the task of Astronaut Photography Localization (APL) with deep learning, combining image retrieval and image matching techniques [35, 5, 36]. Image retrieval is used as a first step, to search within a database of precisely geo-referenced satellite images for those that are most similar to the astronaut photo (called the query). With such a system, a query can be coarsely localized in less than a second, making it a viable solution for such a large scale problem. Afterwards, the retrieval candidates may be refined using more computationally demanding matching techniques. Although these pipelines have shown satisfactory speeds for APL, their performance show a wide margin for improvement, and we hypothesize that this is due to existing APL models being trained only on satellite images [6]. This is a critical limitation, considering that the end goal is to geo-reference astronaut photos, not satellite images.
@@
We hypothesize that it would be beneficial to exploit the available 300,000 hand-labeled open-source astronaut photos to train the image retrieval model. The difficulty with doing so is that these photos are only weakly annotated, i.e., the manual label only provides the geographic location (latitude,longitude) of a random point close to the center of the image. Therefore, these weakly labeled photos cannot be directly used in a metric learning framework, where the model must be provided with pairs of queries (astronaut photos) and database (satellite) images from the same location. In order to meet this criteria, we need these photos to be labeled with their full footprint (i.e. precise GPS location of its four corners), so that we can exactly determine which images from the database can be used as positive samples (i.e., if they have a substantial overlap w.r.t. the training queries). The concept of weak and full labels is visualized in Fig. 2.
@@
Figure 2: Visual example of weak (manual) annotation and full annotation of an astronaut photo. Weak annotation is the geographic coordinates of a single point, which does not provide information about the image’s size (i.e. it could cover a town or an entire continent), whereas full annotation provides coordinates for all 4 corners (called footprint), from which the coordinate of any pixel within the image can be easily calculated (pixel-wise label).
@@
In light of these considerations, we first produce a precise annotation of these 300,000 photos, leveraging their weak annotation and a state-of-the-art matching pipeline to label their full footprint. Then, we demonstrate the importance of using this newly annotated data for training in two ways. First, we create pairs of matching astronaut-satellite images, and use them in a pairwise contrastive loss that takes advantage of the fact that the two images within each pair come from different domains. The creation of these pairs relies on our new annotations, i.e., we select a pair if the two images have enough intersection over union on Earth’s surface. Even with this loss that takes advantage of the newly labeled 300,000 astronaut photos, it is still important to fully leverage the much larger set of all 5.3M geo-referenced satellite images (not just the ones overlapping astronaut photos). However, we observe that the satellite images are uniformly distributed on the surface of the Earth, whereas the photos from astronauts tend to be unevenly distributed, with higher concentrations in salient areas like volcanoes, glaciers and lakes. Therefore, we propose a new mining technique, which allows for a weighted sampling of the large satellite image collection according to the distribution of astronaut photos, which we call Unsupervised Mining, and, to the best of our knowledge, is the first mining technique of its kind. We then make of use of this data with a second contrastive loss.
@@
We find that our training pipeline, yielding a model which we call AstroLoc, leads to a model so powerful that it saturates existing test sets (recall@100 above 99%), and brings us to propose new, more challenging test sets that better reflect the task of APL in the real world. Furthermore, AstroLoc is useful in the real world, and it has already been used to localize hundreds of thousands of images available here [^4] Photos available at: https://eol.jsc.nasa.gov/ExplorePhotos/ . Thanks to AstroLoc, in a few months the backlog of non-localized astronaut photos will be nearly empty for the first time since the launch of the ISS, with the exception of a small fraction that still result in failure cases. Finally, we note that AstroLoc also shows great potential for other applications like ”lost-in-space” satellite orbit determination and historical space image localization.
@@ 2 Related Work
While these solutions have advanced the field of Earth from space localization, they are limited, as the matching-based methods can not achieve the scalability and speed required to localize astronaut photographs in real time. Berton et al. [6] proposes approaching the task as an image retrieval problem, where the database is composed of satellite images of known location. This leads to EarthLoc, the first APL-focused retrieval model. In addition, AnyLoc [20], a universal place recognition model, has been shown to produce strong performance and robustness to the query-database domain gap.
Unmanned aerial vehicle localization A closely related task is that of unmanned aerial vehicle (UAV) localization. Similar to Astronaut Photography Localization (APL), it is approached through cross-domain image retrieval, where the database is often made of satellite imagery and the queries are photos collected by a drone. The goal is to efficiently find the best prediction for a given query in order to localize the drone in real time [42]. Among notable examples from the literature, Bianchi and Barfoot [7] introduced a method utilizing autoencoded satellite images, significantly reducing storage and computation costs while maintaining robust localization performance. Dai et al. [8] proposed an end-to-end framework, FPI, that directly identifies UAV locations by matching UAV-view images with satellite-view counterparts, streamlining the localization process. Li et al. [21] developed a transformer-based adaptive semantic aggregation method, whereas He et al. [16] make use of foundation models to get an estimate of the location to then refine it with visual-inertial odometry algorithms.
@@ 3 Data
Our goal is to localize the extent of astronaut photos (queries) given a worldwide map of geo-referenced satellite images (database). Astronaut photographs represent one of the most diverse and long standing Earth observations data sources. The dataset contains a wide range in spatial resolution, field of view, illumination (including night imagery), and obliquity (astronauts can tilt the camera and take oblique photos). These images have many applications in numerous fields [18, 40, 26, 9, 14, 32, 29], and critically in disaster management and response [33] [^5] https://eol.jsc.nasa.gov/Collections/Disasters/ShowDisastersCollection.pl.
For database images we follow [6] and use a yearly-composited set of cloudless open-source satellite imagery from Sentinel-2, named S2. We use tiles from zoom levels 8-12 [27] (expanding the range of zoom levels used in [6]), to provide a thorough train and test set that reflects the extents of the query images. Characteristics of astronaut and satellite imagery are compared in Tab. 1.
@@ 3.1 Training Set
In contrast to previous work, we propose taking advantage of the 300k manually localized astronaut photos for training. The idea is to pair these photos with matching satellite imagery from the database (i.e. with enough Intersection over Union, or IoU) and apply contrastive training to train models that are robust to the domain differences between astronaut and satellite imagery. Unfortunately, the available manual localization is only a weak annotation (see Fig. 2), which does not provide enough information to compute IoU. To overcome this limitation, we estimate the precise footprint for each query image as follows: for each manually localized query, we select as potential positives all the database images which contain the weak label point at each zoom level, and we rotate these potential positives by 90°, 180° and 270° to maximize the chance of finding a similar image among the potential positives. In practice, given five zoom levels, four rotations, and the fact that each point is covered by four images (there is overlap among the tiles), there are 5×4×4=805$\times$ 4$\times$ 4=805 × 4 × 4 = 80 potential positives per query. Note that different zoom levels are required because it is impossible to estimate an image extent with only a weak label, and a photo could cover an area as small as a city or as big as a continent. We then perform image matching with SuperPoint [10] + LightGlue [22] and the EarthMatch pipeline [5] to get the footprint coordinates of each query.
This method was successful for 221k queries, with the remaining being not localizable due to either errors in the manual label, too much obstruction from clouds, or the image being of the horizon (i.e. Earth limb photograph, with two corners in outer space, hence an invalid footprint). With this now precisely localized collection of astronaut photography, we pair each query with the all database images with an IoU over $t_{iou}=0.2$ producing 865k query-database training pairs. We then discard any pair containing a query that is included within the evaluation sets to avoid any data leakage.
@@ 3.2 Evaluation Sets
We therefore propose new evaluation sets which include all available geolocated queries within an evaluation area, as well as add zoom levels 8 and 12 to the database to match the wider range of query areas. To avoid introducing complexity, we use the same regions as the test sets from [6]. These datasets were named after the geographic location of their center, like Texas, Gobi, and Amazon, so we call these new extended versions Texas-L (L for Large), Gobi-L, etc.
@@ 4.1 Setting
Figure 4: AstroLoc’s training pipeline. The upper branch feeds pairs of matching query-database images to the pairwise loss (Sec. 4.2). The lower branch (Sec. 4.3) first creates clusters of satellite images, then queries are assigned to these clusters, which are sampled according to how many queries are assigned to each cluster (unsupervised mining). A training batch (mined quadruplets, i.e. tuples of 4 images) is sampled from a single cluster and fed to the Multi-similarity Unsupervised Mining (MUM) loss. Queries are not fed to the MUM loss, and are not used to compute the clusters — they are only used to sample training data from a closer distribution to the queries’.
Task Astronaut Photography Localization (APL) is the task of estimating the geographic location covered by a photo (query) taken by astronauts. APL can be approached through image retrieval: for each query, we can estimate its location by finding the most similar match in a large database of geolocated images of Earth.
Formalization Throughout this paper, we rely on three sets of images:
@@
  • The test queries $\mathcal{T}=\{t_{1},t_{2},\ldots,t_{N_{\mathcal{T}}}\}$ is another set of $N_{\mathcal{T}}$ astronaut photos, disjoint from $\mathcal{Q}$, each characterized by its RGB content and timestamp. The timestamp is crucial in determining the location of the ISS when the photo was taken, which enables narrowing down the search space from the entire world surface area (510M sqkm) to the local area visible from the ISS at that time (20M sqkm). Note that due to cosmic radiation bit flips, the timestamp can sometimes be incorrect: in such cases, it is necessary to conduct a worldwide search (see Sec. 5.5).
Training with Astronaut Photos (Queries) While various approaches have been used for APL, no previous work has taken advantage of the huge dataset of 300k astronaut photos labeled by human experts. We posit that their use at training time can lead to sizeable improvements in performance, and we present two novel techniques to use this data. In Sec. 4.2 we show how we use pairs of matching query-database images (i.e., depicting the same location) to implement a pairwise loss (see Fig. 4, upper branch). Additionally, we use the photos a another time, to inform a second loss this time computed on tuples of satellite images that are sampled using the distribution of the queries (see an overview in Fig. 4 bottom). This allows us to use the entire set of database images for training (whereas the pairwise loss can be applied only on areas covered by astronaut photos) while still leveraging the queries (see Sec. 4.3).